For a right triangle with a vertical angle of 40 degrees and horizontal angle of 50 degrees and a horizontal length of 5, what would be the hypotenuse length and vertical length? And if you could show me how you came up with your answer? Thanks in advance!
Answer by alicealc(167)


sin 40 = 5/hypotenuse
0.6428 = 5/hypotenuse
hypotenuse = 5/0.6428 = 7.7785

tan 50 = vertical/5
1.1918 = vertical/5
vertical = 1.1918 * 5 = 5.959